# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v2

@@ 3.3 Shuffle Index
In this subsection we investigate shuffle indices $\chi$. Since DP is inherently concerned with worst-case guarantees, we focus on the upper bound (i.e., the curve $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$) provided by Theorem 3.3. To quantify how tight this upper bound is, we compare it with the lower curve $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$ and examine the width of the resulting band. As $n\to\infty$, this width is asymptotically captured by the ratio $\frac{\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})}{\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})}\;=\;\frac{\underline{\chi}_{\mathrm{lo}}}{\underline{\chi}_{\mathrm{up}}}\,(1+o(1)),$ so that the quantity $\underline{\chi}_{\mathrm{lo}}/\underline{\chi}_{\mathrm{up}}$ provides a convenient measure of the tightness of our blanket-divergence analysis for each local randomizer. We term $\underline{\chi}_{\mathrm{lo}}$ the lower shuffle index and $\underline{\chi}_{\mathrm{up}}$ the upper shuffle index.
Our next result characterizes when $\underline{\chi}_{\mathrm{up}}$ and $\underline{\chi}_{\mathrm{lo}}$ coincide so that the blanket-divergence analysis is asymptotically tight.
Theorem 3.4 (Necessary and sufficient condition for tightness).
@@
Assume the local randomizer $\mathcal{R}$ with blanket mass $\gamma$ satisfies Assumption 2.3. For a pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ define the disagreement set $A(x_{1},x_{1}^{\prime}):=\bigl\{y\in\mathcal{Y}:\mathcal{R}_{x_{1}}(y)\neq\mathcal{R}_{x_{1}^{\prime}}(y)\bigr\},$ that is, the set of outputs where the two local output distributions differ. Assume that the infima in the definitions of $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ are attained by some pairs $(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})$ and $(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}})$, respectively. Then, $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$ if and only if there exists a pair $(x_{1}^{\ast},x_{1}^{\prime\,\ast})$ and an input $x^{\ast}\in\mathcal{X}$ such that
@@
$\chi_{\mathrm{lo}}(x_{1}^{\ast},x_{1}^{\prime\,\ast})=\underline{\chi}_{\mathrm{lo}}\qquad\text{and}\qquad\mathcal{R}_{x^{\ast}}(y)=\gamma\mathcal{R}_{\mathrm{BG}}(y)\ \text{ for almost every }y\in A(x_{1}^{\ast},x_{1}^{\prime\,\ast}).$
@@
In particular, when these equivalent conditions hold, the lower and upper curves in Theorem 3.3 coincide asymptotically, and there exists a unique asymptotic privacy curve $\varepsilon_{n}^{\ast}(\alpha)$ realizing $\delta_{n}=\frac{\alpha}{n}(1+o(1))$ such that $\varepsilon_{n}^{\ast}(\alpha)=\varepsilon_{n}\bigl(\alpha,\underline{\chi}_{\mathrm{lo}}\bigr)\,(1+o(1))$ as $n\to\infty.$
@@
We defer the proof of Theorem 3.4 to Appendix A.4. In particular, the theorem applies to the case of $k$-$R$ ($k\geq 3$), and hence to mechanisms that are built on top of $k$-$R$ or share the same blanket structure, such as local hashing Wang et al. (2017) and (for sufficiently small caps) PrivUnit Asi et al. (2022). Thus, in many practically relevant shuffle protocols, the blanket-divergence analysis is asymptotically optimal in the sense of Theorem 3.4. For more complex families such as generalized Gaussian mechanisms, the upper and lower shuffle indices do not coincide in general, but across a wide range of parameters we empirically observe that their ratio $\underline{\chi}_{\mathrm{lo}}/\underline{\chi}_{\mathrm{up}}$ remains above $0.7$. The detailed calculations of the shuffle indices and their numerical values are deferred to Appendix B.
@@ 4 Analysis of Blanket Divergence via the Fast Fourier Transform (FFT)
The results in the previous section provide an asymptotic characterization of the blanket divergence. However, in applications one is ultimately interested in the finite-$n$ behavior of the shuffled mechanism, and asymptotic expansions alone do not directly yield numerically accurate values of the privacy parameter $\delta$ for concrete choices of $n$.
In this section, we therefore develop a complementary, non-asymptotic approach for computing the blanket divergence at finite $n$. Our goal is to design an algorithm that, for a given local randomizer $\mathcal{R}$ and parameters $(n,\varepsilon)$, produces rigorous upper and lower bounds on $\mathcal{D}^{\mathrm{blanket}}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ with explicitly controlled error and provable running time guarantees. The main idea is to exploit the same structure that underlies our CLT-based analysis: the blanket divergence can be expressed in terms of sums of i.i.d. privacy amplification random variables. This observation allows us to approximate the distribution of the sum by means of the Fast Fourier Transform (FFT). To obtain rigorous bounds rather than heuristic approximations, we carefully track the errors introduced by truncation of the summands, discretization on a grid, and FFT wrap-around. By combining these error bounds with the asymptotic expansion from Lemma 3.2, we can both tune the numerical parameters (window size, grid resolution, truncation level) and analyze the overall complexity of the method.
@@
Remark 4.1 (Transformation of the blanket divergence for FFT).
Let $M\sim 1+\mathrm{Bin}(n-1,\gamma)$. Then, $\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\left[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\right]-\mathrm{e}^{\varepsilon}\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\left[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\right].$
The proof is provided in Appendix A.5. Compared with the transformations in Lemma 3.1, here we work with a representation of the blanket divergence that does not involve explicit expectations or integrals, which is particularly convenient for deriving rigorous bounds on the truncation error. In particular, in our FFT-based computation we must truncate the support of the random variable in order to approximate an a priori unbounded distribution on a finite grid. This issue is especially relevant for non-pure local DP (see, e.g., the discussion of the Gaussian mechanism in Balle et al. Balle et al. (2019)); hence this formulation is crucial for our purposes.
Su et al. Su et al. (2025) also propose an FFT-based accountant for the blanket divergence, but their running time scales as $O(n^{2})$ in the number of users when targeting a constant additive error. In contrast, by adopting the concentration-inequality-based error control of Gopi et al. Gopi et al. (2021) with our results in Section 3, we can tune the truncation and discretization parameters so that our algorithm runs in near-linear time $\widetilde{O}(n)$ while achieving a prescribed relative error. Algorithms 1 and 2 together form our FFT-based procedure.
@@ 4.1 FFT Algorithm
Theorem 4.2 (Error bounds for FFT-based blanket divergence computation).
@@
Fix a distinct pair $x_{1},x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}$ be a local randomizer. Let the reference distribution be $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ For $Y\sim\mathcal{R}_{\mathrm{ref}}$, denote by $F_{\mathrm{ref}}$ the CDF of the privacy amplification random variable $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$, and for $j\in\{x_{1},x_{1}^{\prime}\}$ denote by $F_{j}$ the CDF of $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ under $Y\sim\mathcal{R}_{j}$. Let $(B_{i})_{i=2}^{n}$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$ and $(Y_{i})_{i=2}^{n}$ be i.i.d. with common distribution $\mathcal{R}_{\mathrm{ref}}$, independent of $(B_{i})_{i=2}^{n}$.
Given parameters $(n,\gamma,F_{\mathrm{ref}},s,w^{\mathrm{in}},h,w^{\mathrm{out}})$, we define the truncation event $\perp:=\Bigl\{\exists i\in\{2,\dots,n\}:B_{i}=1,\ l_{\varepsilon}\bigl(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}\bigr)\notin[s,s+w^{\mathrm{in}}]\Bigr\}$ and let $\bigl(\mathbf{p}_{S-\mu_{S}^{\mathrm{di}}},\mu_{S}^{\mathrm{tr}}\bigr):=\textsc{CalculatePMF}\bigl(n,\gamma,F_{\mathrm{ref}},s,w^{\mathrm{in}},h,w^{\mathrm{out}}\bigr).$ For $x\in\{x_{1},x_{1}^{\prime}\}$ and $c\geq 0$, define the main probability term
$P(c;x):=\bigl(1-\Pr[\perp]\bigr)\,\textsc{CalculateMainTerm}\bigl(\mathbf{p}_{S-\mu_{S}^{\mathrm{di}}},\mu_{S}^{\mathrm{tr}},F_{x},c\bigr).$

### v3

@@ 3.4 Shuffle Index
In this subsection, we explore the shuffle index $\chi$: its interpretation, examples, and the analysis of the tightness of the blanket-divergence bounds through the shuffle index.
Interpretation of Shuffle Index.
Technically, by the CLT as shown in Section 3.2, the leading asymptotics of the blanket divergence with $(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ can be summarized by the two parameters $\gamma$ and $\sigma$. Here we defined $\sigma^{2}:=\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{ref}}}\left(\frac{\mathcal{R}_{x_{1}}(Y)-\mathcal{R}_{x_{1}^{\prime}}(Y)}{\mathcal{R}_{\mathrm{ref}}(Y)}\right)=\int_{\mathcal{Y}}\frac{\bigl(\mathcal{R}_{x_{1}}(y)-\mathcal{R}_{x_{1}^{\prime}}(y)\bigr)^{2}}{\mathcal{R}_{\mathrm{ref}}(y)}\,dy$ so that $\sigma$ is a $\chi^{2}$-type distance between $\mathcal{R}_{x_{1}}$ and $\mathcal{R}_{x_{1}^{\prime}}$ with respect to the reference distribution $\mathcal{R}_{\mathrm{ref}}$. This differs from the usual $\chi^{2}$ divergence, whose reference distribution is $\mathcal{R}_{x_{1}}$ or $\mathcal{R}_{x_{1}^{\prime}}$; here the reference distribution is the third distribution $\mathcal{R}_{\mathrm{ref}}$. In Fig. 3, we take $\mathcal{R}$ to be Gaussian and plot the integrand of $\sigma^{2}$, namely $(\mathcal{R}_{x_{1}}-\mathcal{R}_{x_{1}^{\prime}})^{2}/\mathcal{R}_{\mathrm{BG}}$. Compared to the ordinary squared distance $(\mathcal{R}_{x_{1}}-\mathcal{R}_{x_{1}^{\prime}})^{2}$, we see that regions where the blanket is thin are emphasized. This distinction represents the shuffled setting: we can interpret $\sigma$ as distinguishability of $\mathcal{R}_{x_{1}}$ and $\mathcal{R}_{x_{1}^{\prime}}$ from the blanket $\mathcal{R}_{\mathrm{BG}}$ view.
To reiterate, the blanket divergence can be interpreted as the natural upper bound of the hockey-stick divergence between $\mathcal{S}(\mathcal{R}_{x_{1}},\mathcal{R}_{\mathrm{ref}},\ldots,\mathcal{R}_{\mathrm{ref}})$ and $\mathcal{S}(\mathcal{R}_{x_{1}^{\prime}},\mathcal{R}_{\mathrm{ref}},\ldots,\mathcal{R}_{\mathrm{ref}})$. $\gamma$ represents the fraction of messages that fall into the blanket, and hence the effective message size is $M\approx n\gamma$. On the other hand, since $1/\sigma$ is the reciprocal of the distinguishability measure in the above setting, it quantifies the per-message indistinguishability. Because the contribution of $\gamma$ to indistinguishability appears through the $\sqrt{n}$ scaling of fluctuations, the index aggregates these effects as $\sqrt{\gamma}/\sigma$.
Examples of Shuffle Indices
Table 1: Shuffle index asymptotics of $\beta$-Gauss with $\mathrm{Var}(\mathcal{R}(x))=\sigma_{0}^{2}$ where $c_{\beta}=\frac{\Gamma(1/\beta)}{\beta\sqrt{\Gamma(3/\beta)\Gamma(2-1/\beta)}}$ and $k$-$R$.
$\beta$-Gauss $\sigma_{0}\to\infty$ $\sigma_{0}\to 0$ $\chi_{\mathrm{lo}}$ $c_{\beta}\sigma_{0}\left(1+o\left(1\right)\right)$ $\exp\!\bigl(-\Omega(\sigma_{0}^{-\beta})\bigr)$ $\chi_{\mathrm{lo}}/\chi_{\mathrm{up}}$ $1-O\!\left(\frac{1}{\sigma_{0}}\right)$ $\frac{1}{\sqrt{2}}+o(1)$ kkRR $\mid$ 22RR $\varepsilon_{0}\to 0$ $\varepsilon_{0}\to\infty$ $\chi_{\mathrm{lo}}$ $\sqrt{\frac{k}{2}}\frac{1}{\varepsilon_{0}}\left(1+O(\varepsilon_{0})\right)$ $\frac{e^{-\frac{\varepsilon_{0}}{2}}}{\sqrt{2}}\left(1+O(e^{-\varepsilon_{0}})\right)$ $\chi_{\mathrm{lo}}/\chi_{\mathrm{up}}$ 11 $\mid$ $O(\varepsilon_{0}) 11 \mid$ $\frac{1}{\sqrt{2}}+O(e^{-\varepsilon_{0}})$
We illustrate the shuffle index of two local randomizers: the $\beta$-Gaussian mechanism and $k$-$R$. Concretely, we examine (i) the asymptotics of the lower shuffle index $\chi_{\mathrm{lo}}$, and (ii) the ratio $\chi_{\mathrm{lo}}/\chi_{\mathrm{up}}$ as a proxy for the tightness of the blanket-divergence sandwich. Closed-form derivations and some plots of shuffle indices are deferred to Appendix C. The resulting asymptotics are summarized in Table 1: for $\beta$-Gaussian we parametrize by the local noise scale $\sigma_{0}$, while for $k$-$R$ we parametrize by the pure local-DP level $\varepsilon_{0}$, and in both cases we report limits as the parameter tends to 0 and $\infty$. These asymptotics precisely capture the fundamental behavior established in theoretical bounds Erlingsson et al. (2019) and explain recent empirical observations Chen et al. (2023) that shuffling yields highly efficient privacy amplification in the high-privacy regime, while providing negligible amplification in the low-privacy regime.
$\beta$-Gaussian. For large noise ($\sigma_{0}\to\infty$), the lower shuffle index grows linearly, $\chi_{\mathrm{lo}}\approx c_{\beta}\sigma_{0}$, where $c_{\beta}$ depends only on the shape parameter $\beta$. Concretely, $c_{\beta}$ increases from $1/\sqrt{2}$ at $\beta=1$ (Laplace) up to $1$ at $\beta=2$ (Gaussian), and then decreases for $\beta>2$; hence, in the high-noise regime, the Gaussian case $\beta=2$ yields the largest shuffle index. Moreover, the tightness ratio converges to $1$, so the blanket-divergence bounds essentially coincide asymptotically when $\sigma_{0}$ is sufficiently large. That is, in the mean estimation task, since $\sigma_{0}$ monotonically determines the accuracy of the estimate, $\beta=2$ induces the best privacy-utility trade-off in the high-noise regime (see Section 5.3 for the exploration of this topic). In contrast, in the low-noise regime ($\sigma_{0}\to 0$), $\chi_{\mathrm{lo}}$ decays exponentially to 0, which reflects that shuffling provides little amplification when the local noise is too small.
Remark 3.5 (Dimensional dependence).
For the Gaussian case ($\beta=2$), it is natural to extend the input domain to $d$ dimensions. Concretely, consider the Gaussian local randomizer $\mathcal{R}(x)=x+\mathcal{N}(0,\sigma_{0}^{2}I_{d})$ with fixed $\sigma_{0}$ and inputs in an $\ell_{2}$-ball of fixed radius. Then the blanket mass satisfies $\gamma=\exp(-\Theta(\sqrt{d}))$, i.e., it decays exponentially fast in $\sqrt{d}$. Intuitively, this decay is driven by the thin-shell phenomenon of high-dimensional Gaussians: a typical output exponentially concentrates on a thin shell around its mean $x$ (i.e., $\|\mathcal{R}(x)-x\|\approx\sigma_{0}\sqrt{d}$). On this typical shell, the blanket distribution takes an infimum over all inputs in the $\ell_{2}$-ball, so the blanket becomes exponentially thin in $\sqrt{d}$ (see Section C.1.4 for details). Moreover, while $\chi_{\mathrm{up}}$ is not directly affected by $\gamma$, $\chi_{\mathrm{lo}}$ is suppressed by the factor $\sqrt{\gamma}$ and thus shrinks exponentially with $\sqrt{d}$; consequently, the gap between the corresponding upper and lower bounds becomes large. Closing this gap appears to require techniques that go beyond the blanket divergence, which remains an open problem.
$k$-$R$. In large noise ($\varepsilon_{0}\to 0$), we have $\chi_{\mathrm{lo}}\approx\sqrt{\frac{k}{2}}\frac{1}{\varepsilon_{0}}$. Importantly, for $k\geq 3$ the tightness ratio satisfies $\chi_{\mathrm{lo}}/\chi_{\mathrm{up}}=1$, i.e., the two indices coincide and the blanket-divergence sandwich is tight at leading order; for $2$-$R$, the ratio is also close to $1$ as $\varepsilon_{0}\to 0$. In the low-noise regime ($\varepsilon_{0}\to\infty$), $\chi_{\mathrm{lo}}$ again vanishes exponentially, and thus shuffling becomes ineffective.
Remark 3.6 (Comparison to the clone paradigm).
As noted in Remark 3.1, the clone paradigm can be understood in terms of a blanket divergence with $\gamma=1$ and an appropriate choice of the categorical distributions. Hence, we can induce a shuffle index of the clone paradigm. For $k$-$R$, the explicit reduction of Feldman et al. (2023, Theorem 7.4) leads to a shuffle index that coincides with our $\chi_{\mathrm{lo}}$ (see Section C.2.3 for details). While their work only demonstrates empirical closeness to known lower bounds, our framework provides a theoretical and asymptotic explanation for the closeness.
Tightness of Blanket-Divergence Analysis.
We have seen that $\chi_{\mathrm{lo}}/\chi_{\mathrm{up}}=1$ for $k$-$R$ with $k\geq 3$. Here, we explore the structural conditions under which this equality holds, which implies that the blanket-divergence bounds are asymptotically tight in the sense of Theorem 3.4. Formally, we have the following characterization.
Theorem 3.7.
@@
Let $\mathcal{R}\in\mathfrak{R}$ with blanket mass $\gamma$ and assume that the suprema in the definitions of $\chi_{\mathrm{lo}}(\mathcal{R})$ and $\chi_{\mathrm{up}}(\mathcal{R})$ are attained by some pairs. Then, $\chi_{\mathrm{up}}(\mathcal{R})=\chi_{\mathrm{lo}}(\mathcal{R})$ if and only if there exists a pair which attains the suprema for the lower shuffle index $(x_{1}^{\ast},x_{1}^{\prime\,\ast})$ and an input $x^{\ast}\in\mathcal{X}$ such that
@@
$\mathcal{R}_{x^{\ast}}(y)=\gamma\mathcal{R}_{\mathrm{BG}}(y)\ \text{ for almost every }y\in\bigl\{y\in\mathcal{Y}:\mathcal{R}_{x_{1}^{\ast}}(y)\neq\mathcal{R}_{x_{1}^{\prime\,\ast}}(y)\bigr\}.$
@@
In particular, when these equivalent conditions hold, the lower and upper curves in Theorem 3.4 coincide asymptotically, and there exists an asymptotic privacy curve $\varepsilon_{n}^{\ast}(\alpha)$ realizing $\delta_{\mathcal{S}\circ\mathcal{R}^{n}}(\varepsilon_{n}^{\ast}(\alpha))=\frac{\alpha}{n}(1+o(1))$ such that $\varepsilon_{n}^{\ast}(\alpha)=\varepsilon_{n}\bigl(\alpha,\chi_{\mathrm{lo}}(\mathcal{R})\bigr)\,(1+o(1))$ as $n\to\infty.$
@@
We defer the proof of Theorem 3.7 to Appendix A.3. In particular, the theorem applies to the case of $k$-$R$ ($k\geq 3$), and hence to mechanisms that are built on top of $k$-$R$ or share the same blanket structure, such as local hashing Wang et al. (2017) and (for sufficiently small caps) PrivUnit Asi et al. (2022). Thus, in many practically relevant shuffle protocols, the blanket-divergence analysis is asymptotically optimal in the sense of Theorem 3.7.
@@ 4 Analysis of Blanket Divergence via the Fast Fourier Transform (FFT)
The results in the previous section provide an asymptotic characterization of the blanket divergence. However, in applications one is ultimately interested in the finite-$n$ behavior of the shuffled mechanism. To this end, we develop a complementary, non-asymptotic approach for computing the blanket divergence at finite $n$.
Our goal is to design an algorithm that, for a given local randomizer $\mathcal{R}$ and parameters $(n,\varepsilon)$, produces rigorous upper and lower bounds on $\mathcal{D}^{\mathrm{blanket}}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ with explicitly controlled error and provable running time guarantees. The main idea is to exploit the same structure that underlies our CLT-based analysis: the blanket divergence can be expressed in terms of sums of i.i.d. privacy amplification random variables. This observation allows us to approximate the distribution of the sum by means of the Fast Fourier Transform (FFT). To obtain rigorous bounds rather than heuristic approximations, we carefully track the errors introduced by truncation of the summands, discretization on a grid, and FFT wrap-around. By combining these error bounds with the asymptotic expansion from Lemma 3.2, we can both tune the numerical parameters (window size, grid resolution, truncation level) and analyze the overall complexity of the method.
@@
Remark 4.1.
Let $M\sim 1+\mathrm{Bin}(n-1,\gamma)$. Then,
$\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\mathrm{blanket}}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\left[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\right]-\mathrm{e}^{\varepsilon}\Pr_{\begin{subarray}{c}Y_{1}\sim\mathcal{R}_{x_{1}^{\prime}}\\ Y_{2:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\left[\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\right].$
The proof is provided in Appendix A.4. Here, we work with a representation of the blanket divergence that does not involve explicit expectations or integrals, which is particularly convenient for deriving rigorous bounds on the truncation error. In particular, in our FFT-based computation we must truncate the support of the random variable in order to approximate an unbounded distribution on a finite grid. This issue is relevant for non-pure local DP (see, e.g., the discussion of the Gaussian mechanism in Balle et al. Balle et al. (2019)); hence this formulation is crucial for our purposes.
Su et al. Su et al. (2025) also propose an FFT-based accountant for the blanket divergence, but their running time scales as $O(n^{2})$ in the number of users when targeting a constant additive error. In contrast, by adopting the concentration-inequality-based error control of Gopi et al. Gopi et al. (2021) with our results in Section 3, we can tune the truncation and discretization parameters so that our algorithm runs in near-linear time $\widetilde{O}(n)$ while achieving a prescribed relative error.
@@ 4.1 FFT Algorithm
Theorem 4.2.
@@
Fix a distinct pair $x_{1},x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}$ be a local randomizer. Let the reference distribution be $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ For $Y\sim\mathcal{R}_{\mathrm{ref}}$, denote by $F_{\mathrm{ref}}$ the CDF of the privacy amplification random variable $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$, and for $j\in\{x_{1},x_{1}^{\prime}\}$ denote by $F_{j}$ the CDF of $l_{\varepsilon}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$ under $Y\sim\mathcal{R}_{j}$. Let $(B_{i})_{i=2}^{n}$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$ and $(Y_{i})_{i=2}^{n}$ be i.i.d. with $\mathcal{R}_{\mathrm{ref}}$, independent of $(B_{i})_{i=2}^{n}$.
Given parameters $(n,\gamma,F_{\mathrm{ref}},s,w^{\mathrm{in}},h,w^{\mathrm{out}})$, we define the truncation event $\perp:=\Bigl\{\exists i\in\{2,\dots,n\}:B_{i}=1,\ l_{\varepsilon}\bigl(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}}\bigr)\notin[s,s+w^{\mathrm{in}}]\Bigr\}$ and let $\bigl(\mathbf{p}_{S-\mu_{S}^{\mathrm{di}}},\mu_{S}^{\mathrm{tr}}\bigr):=\textsc{CalculatePMF}\bigl(n,\gamma,F_{\mathrm{ref}},s,w^{\mathrm{in}},h,w^{\mathrm{out}}\bigr).$ For $x\in\{x_{1},x_{1}^{\prime}\}$ and $c\geq 0$, define $P(c;x):=\bigl(1-\Pr[\perp]\bigr)\,\textsc{CalculateMainTerm}\bigl(\mathbf{p}_{S-\mu_{S}^{\mathrm{di}}},\mu_{S}^{\mathrm{tr}},F_{x},c\bigr).$